OpenStudy (anonymous):

What number must you add to complete the square? x2 + 4x = 12

OpenStudy (anonymous):

the x2 = x^2 sorry

OpenStudy (anonymous):

you have to add 4 to both sides. then equation can be solved

OpenStudy (anonymous):

x^2+2(2)(x)+2^2 - 2^2=12

OpenStudy (anonymous):

what is (x+2)^2 =?